A cube has a surface area of 240 cm^2 . What is the length of one side?
Answer by mananth(16946)

L*L = l^2 = surface area of 1 side
there are 6 sides to a cube
6L^2= 240
divide by 6
6L^2/6 = 240/6
L^2= 40
L = sqrt 40